OpenStudy (sodapop):

Using these similarities and differences, how would you transform f(x) = 2 sin(2x - π) + 3 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (anonymous):

sin(a-pi)= cos (pi/2 -a+pi)=cos (3pi/2-a)=cos(a- 3 pi/2) so here, 2 sin(2x-pi)+3= 2 cos(2x- 3 pi/2) +3

OpenStudy (sodapop):

Thank you